# Non-Linear Model-Based Sequential Decision-Making in Agriculture

### v1

@@ Abstract
Sequential decision-making is central to sustainable agricultural management and precision agriculture, where resource inputs must be optimized under uncertainty and over time. However, such decisions must often be made with limited observations, whereas classical bandit and reinforcement learning approaches typically rely on either linear or black-box reward models that may misrepresent domain knowledge or require large amounts of data. We propose a family of nonlinear, model-based bandit algorithms that embed domain-specific response curves directly into the exploration-exploitation loop. By coupling (i) principled uncertainty quantification with (ii) closed-form or rapidly computable profit optima, these algorithms achieve sublinear regret and near-optimal sample complexity while preserving interpretability. Theoretical analysis establishes regret and sample complexity bounds, and extensive simulations emulating real-world fertilizer-rate decisions show consistent improvements over both linear and nonparametric baselines (such as linear UCB and $k$-NN UCB) in the low-sample regime, under both well-specified and shape-compatible misspecified models. Because our approach leverages mechanistic insight rather than large data volumes, it is especially suited to resource-constrained settings, supporting sustainable, inclusive, and transparent sequential decision-making across agriculture, environmental management, and allied applications. This methodology directly contributes to SDG 2 (Zero Hunger) and SDG 12 (Responsible Consumption and Production) by enabling data-driven, less wasteful agricultural practices.
@@ 1 Introduction
Identifying the most effective crop management practice for a given site is a recurring challenge in agriculture. While field trials and expert recommendations provide useful starting points, the actual performance of management strategies such as different fertilizer rates, planting densities, or irrigation schedules can be highly variable due to differences in soil properties, previous management, and other biological or environmental factors (Hochman and Carberry,, 2011; Tilman et al.,, 2002). As a result, farmers and researchers must repeatedly decide which management actions to implement each season, often with limited prior information about their true effects. This problem exemplifies sequential decision-making under uncertainty, where the goal is to efficiently learn and optimize management choices over time.
@@
In the statistical and machine learning literature, such problems are naturally framed as multi-armed bandit problems, dating back to Robbins, (1952). They have been widely applied in areas such as precision medicine (Lu et al.,, 2021), recommendation systems (Li et al.,, 2010), and, increasingly, precision agriculture (Gautron et al.,, 2022). The term originates from the analogy of a gambler choosing between several slot machines (“one-armed bandits”), each with unknown payout rates. At each round, the decision-maker selects an action/arm (e.g., a specific management practice) and observes its outcome (e.g., crop yield or profit). The central challenge is to balance exploration (trying new or uncertain options to learn about their effects) with exploitation (repeating choices that have performed well so far) in order to maximize cumulative reward or minimize loss over repeated decisions. Bandit algorithms provide principled, data-driven strategies to manage this tradeoff, making them well-suited to resource optimization tasks in agriculture, environmental management, and beyond. However, most existing bandit methods assume linear, generalized linear, or black-box (non-parametric) reward models that often fail to capture the mechanistic, nonlinear dose-response dynamics common in these domains.
@@
In this work, we develop and analyze a new family of bandit algorithms that directly incorporate domain-specific nonlinear response models, offering greater interpretability and sample efficiency compared to classical linear or black-box approaches. Our approach is particularly valuable in settings where repeated decision-making is required but available data is limited and the underlying input-response relationship is nonlinear or complex.
Nonlinear models can encompass a wide range of mean functions. Although individual nonlinear models may be less flexible than general linear models (e.g., polynomials), they are often more parsimonious, interpretable, and tailored to domain-specific processes. Their interpretability stems from the ability to associate parameters with biologically meaningful phenomena. In agriculture, nonlinear models (Miguez et al.,, 2018) refer to mathematical equations that describe complex relationships between variables, such as crop yield response to fertilizer, temperature effects on plant growth, or pest spread–where the response is not linear but follows more intricate patterns (e.g., exponential or logistic). The adoption of such models is motivated not only by statistical fit but also by the need for transparent, explainable recommendations in sustainable resource management. Let us first provide a brief overview of research developments in sequential decision-making suitable for our discussion and motivation for developing new methodology and algorithms.
@@ 1.1 Background and related work
Multi-armed bandit algorithms (Berry and Fristedt,, 1985; Lattimore and Szepesvári,, 2018) provide a core statistical framework for sequential decision-making under uncertainty, where the key challenge is to balance exploration and exploitation. While the original bandit formulation assumed independent, identically distributed (i.i.d.) rewards, subsequent work has extended the framework to incorporate contextual information leading to the development of contextual bandit models (Slivkins,, 2014; Chu et al.,, 2011). In these settings, regression models such as linear (Abbasi-Yadkori et al.,, 2011), generalized linear models (GLMs) (Li et al.,, 2010; Filippi et al.,, 2010), and more flexible nonparametric approaches, including kernels, nearest neighbors, and neural networks (Yang and Zhu,, 2002; Rigollet and Zeevi,, 2010; Srinivas et al.,, 2010; Valko et al.,, 2013) enable data-driven, adaptive decision-making tailored to evolving arm/covariate information.
Classical nonlinear regression models, such as Mitscherlich (Dhanoa et al.,, 2022), Michaelis-Menten (Lopez et al.,, 2000), quadratic threshold models (Bélanger et al.,, 2000) and logistic response functions (Sepaskhah et al.,, 2011), are routinely used to capture crop yield response to fertilizer and other management variables. These models offer interpretable parameters tied to underlying biological processes, which is important for transparent and actionable recommendations (De Bruin and Pedersen,, 2008; Miguez et al.,, 2018).
@@
It is well-known that model-based bandit and reinforcement learning approaches can be far more sample-efficient than model-free methods when the underlying reward mechanisms are known (Osband and Van Roy,, 2014; Sun et al.,, 2019). However, such approaches have received limited attention in agricultural decision-making from a precision-agriculture and sequential decision-making perspective, where established nonlinear models are often available but underutilized.
Our work addresses this gap by integrating domain-driven nonlinear models directly into the bandit framework. This enables adaptive, interpretable, and sample-efficient resource management, and advances the practical and theoretical toolkit for sustainable agriculture and related applications.
@@
  • Theoretical guarantees: We provide regret and sample complexity results under both well-specified and shape-compatible misspecified models, and discussing the implications in the agricultural sequential decision-making domain.
  • Simulation studies emulating real-world practice: Through extensive simulations inspired by real-world fertilizer optimization under both well-specified and misspecified settings, we demonstrate the practical benefits of our approach over classical linear and nonparametric bandit methods, particularly in low-sample regimes.
  • Interpretability and domain relevance: By leveraging mechanistic models with interpretable parameters, our methods yield actionable recommendations for practitioners, supporting sustainable, transparent, and inclusive decision-making.
We begin by setting up the problem in Section 2. In Section 3, we lay out the proposed framework for nonlinear model-based bandits, in which we study three model-based algorithms: $\epsilon$-greedy, Upper Confidence Bound (UCB), and the ViOlin algorithm. Section 4 presents known results for parametric contextual bandits in both well-specified and misspecified settings, and derives sample complexity bounds for the specific nonlinear models considered by the ViOlin algorithm. We then conduct extensive simulation experiments emulating real-world yield production and fertilizer optimization scenarios, with the goal of profit maximization, especially in sample-limited settings. Finally, Section 6 concludes the paper.
2 Problem setup: Model-based nonlinear bandits
In agricultural systems, sequential decisions regarding inputs such as fertilizer rates, irrigation levels, or planting densities must be made under uncertainty, often with limited experimental data. These decisions involve selecting from a continuous (or discrete) set $\mathcal{X}$ of input levels (or arms) based on partially known relationships between inputs and outcomes, such as crop yield or profit. This setting naturally corresponds to a multi-armed bandit problem, where the goal is to sequentially identify input levels that maximize cumulative outcomes over time.
Formally, we consider a sequential decision-making scenario with a time horizon of $T$ rounds. In practice, each round can represent a growing season or field trial, with decisions made at the start and outcomes observed at harvest. At each round $t=1,\dots,T$, the decision-maker (agent) chooses an action (input level) $x_{t}\in\mathcal{X}$ from a predefined set of feasible treatments, such as nitrogen application rates. After choosing action $x_{t}$, the agent observes a noisy reward $y_{t}$, representing yield or profit, modeled as:
@@ 2 Problem setup: Model-based nonlinear bandits
This assumption ensures controlled tail behavior of the noise distribution, supporting reliable statistical concentration results.
The decision-maker seeks a policy/algorithm $\pi$ that selects actions at any given round based on historical observations to maximize cumulative expected reward. Equivalently, performance can be assessed through the cumulative regret, which quantifies the total expected loss incurred by not always selecting the optimal arm at each round. Specifically, the cumulative regret after $T$ rounds is defined as:
@@
where $x_{t}^{*}=\arg\max_{x\in\mathcal{X}}f(x;\beta^{*})$ denotes the optimal treatment at round $t$, which is unknown to the decision-maker. Note that the cumulative regret $R_{T}(\pi)$ is a random quantity since it depends on the actions chosen by the policy $\pi$, which are themselves functions of the observed data.
The objective is to devise algorithms achieving sublinear regret, meaning $R_{T}(\pi)=o_{P}(T)$, so that the average regret per round diminishes over time. Such a property guarantees that the policy rapidly identifies near-optimal input levels, minimizing wasted resources and experimental costs. Other performance metrics more suitable for risk-averse decision-making have also been proposed, such as the Conditional Value-at-Risk (CVaR) of the regret, which emphasizes control of the worst-case outcomes in the tail of the regret distribution (Baudry et al.,, 2021; Wang et al.,, 2023).
To illustrate our methodology, we focus on fertilizer-yield relationships and specifically investigate four biologically motivated nonlinear models commonly employed in agricultural research:
@@
    Michaelis-Menten Model: $f(x;a,b,d)=d+\frac{ax}{b+x}$, describing saturating nutrient uptake or growth responses.
@@
    Quadratic Threshold Model:
@@
    modeling symmetric or asymmetric yield responses that are often observed empirically, often with a threshold term to denote a plateau.
@@
    Logistic Dose-Response Model: $f(x;A,B,C,d)=d+\frac{A}{1+\exp\left(-B(x-C)\right)}$, frequently used to model threshold or inflection point responses to inputs.
It is important to emphasize here that these models offer interpretable parameters tied to agronomic phenomena (e.g., saturation, growth rate, efficiency), making them not only statistically efficient but also scientifically meaningful. For example, in the Mitscherlich model, where $A$ denotes the maximum additional yield achievable by the input, $b$ controls the rate of response or efficiency of the input, and $d$ is the baseline yield without the input. A larger value of $b$ indicates that the yield saturates more rapidly with increasing input, whereas a smaller $b$ corresponds to a more gradual response. These parameters are directly tied to agronomic processes, making the resulting policy recommendations both transparent and actionable for practitioners. Beyond these specific models, the proposed framework is broadly applicable to agricultural management problems. These include models for biomass accumulation over time, plant growth as a function of temperature, light-use efficiency, and nitrogen loss in response to rainfall, many of which are well captured by classical nonlinear models. A comprehensive list of such models and their applications is given by Miguez et al., (2018) and references therein. Detailed descriptions, parameter choices, and practical implementations of these models are presented explicitly in simulation studies (Section 5).
3 Algorithms for model-based nonlinear contextual bandits
Bandit algorithms address the fundamental challenge of sequential decision-making under uncertainty, where one must choose among several options (arms) over time to maximize cumulative reward. Recall, these algorithms are typically designed to achieve two goals: (1) balancing the exploration-exploitation trade-off, and (2) minimizing cumulative regret. In this section, we describe core algorithmic strategies for nonlinear contextual bandit problems, where the expected reward is modeled as a nonlinear function of the arms and, possibly, covariates. Below, we provide a general algorithmic template, followed by detailed descriptions of three model-based strategies, namely, $\epsilon$-greedy, Upper Confidence Bound (UCB), and ViOlin, highlighting how each balances exploration and exploitation. Note that all the three algorithms have been studied in various bandit problems before (Auer and Ortner,, 2010; Arya and Yang,, 2020; Dong et al.,, 2021), we propose model-based versions here.
@@ 3.1 General framework
Algorithm 1 General Model-based Nonlinear Contextual Bandit Framework
@@
where $\ell$ is a loss function (such as squared-error loss) computed using the observed $(x_{i},y_{i})$ pairs, with the goal of minimizing the loss. This practical approach is central to our implementation, allowing bandit algorithms to leverage data-driven parameter estimates and domain-specific models. Then, in step 5, the decision-maker chooses the next action by balancing the exploration-exploitation trade-off and satisfying the criterion for cumulative reward maximization. Then, in step 6, the reward is observed after the season and is added to the dataset.
@@
Unlike standard supervised learning or a fully-factorial experiments, which provides full feedback by revealing the outcome for every possible action/arm, bandit algorithms operate under partial feedback: at each round, only the reward corresponding to the action actually chosen is observed. This is often referred to as bandit feedback.
@@ 3.2 Epsilon-Greedy
The $\epsilon$-greedy algorithm (Algorithm 2) (Langford and Zhang,, 2008; Yang and Zhu,, 2002) is one of the simplest approaches for balancing exploration and exploitation. As the name suggests, most of the time it is greedy, i.e., it chooses the action that appears best according to current knowledge (exploitation), but with small probability $\epsilon$, it selects an action at random (exploration). This mechanism ensures continued learning, preventing the algorithm from becoming stuck on suboptimal choices.
@@
Algorithm 2 Epsilon-Greedy for Model-based Nonlinear Contextual Bandits

### v2

# Nonlinear Model-Based Sequential Decision-Making in Agriculture

@@ Abstract
Agricultural decision-making faces a dual challenge: sustaining high yields to meet global food security needs while reducing the environmental impacts of fertilizer use. Nitrogen inputs are central to this tension. They are indispensable for crop growth yet also major drivers of greenhouse gas emissions, nutrient runoff, and escalating production costs. Addressing these intertwined pressures requires adaptive decision-support tools that are not only statistically principled but also economically sustainable and interpretable for practitioners.
We develop nonlinear model-based bandit algorithms as a framework for adaptive fertilizer management under uncertainty. Building on classical mechanistic yield-response models including Mitscherlich, Michaelis-Menten, quadratic plateau, and logistic functions, our approach links algorithmic exploration-exploitation strategies directly to interpretable biological processes such as maximum yield potential and nutrient efficiency. This grounding makes recommendations transparent for practitioners while supporting cost-effective and sustainable input use. Methodologically, we establish regret and sample complexity results for the well-specified nonlinear case, examine robustness under misspecification, and conduct extensive simulations reflecting realistic agricultural scenarios. The results reveal a key trade-off: nonparametric methods reduce short-term variability but introduce persistent bias, whereas model-based approaches capture biological structure and deliver more reliable long-term profit, ensuring economically sustainable fertilizer use. Our findings highlight how interpretable, uncertainty-aware algorithms can strengthen farmer livelihoods and contribute to more efficient and sustainable use of agricultural inputs.
@@ 1 Introduction
Identifying the most effective crop management practice for a given site is a recurring challenge in agriculture. While field trials and expert recommendations provide useful starting points, the actual performance of management strategies such as different fertilizer rates, planting densities, or irrigation schedules can be highly variable due to differences in soil properties, previous management, and other biological or environmental factors (Hochman and Carberry,, 2011; Tilman et al.,, 2002). As a result, farmers and researchers must repeatedly decide which management actions to implement each season, often with limited prior information about their true effects. This problem exemplifies sequential decision-making under uncertainty, where the goal is to efficiently learn and optimize management choices over time in a way that supports both productivity and sustainability.
@@
In the statistical and machine learning literature, such problems are naturally framed as multi-armed bandit problems, dating back to Robbins, (1952). They have been widely applied in areas such as precision medicine (Lu et al.,, 2021), recommendation systems (Li et al.,, 2010), and, increasingly, agriculture (Gautron et al.,, 2022). The term originates from the analogy of a gambler choosing between several slot machines (“one-armed bandits”), each with unknown payout rates. At each round, the decision-maker selects an action/arm (e.g., a specific management practice) and observes its outcome (e.g., crop yield or profit). The central challenge is to balance exploration (trying new or uncertain options to learn about their effects) with exploitation (repeating choices that have performed well so far). Bandit algorithms provide principled, data-driven strategies to manage this tradeoff, making them well-suited to resource optimization tasks in agriculture and environmental management. However, most existing bandit methods assume linear, generalized linear, or black-box (non-parametric) reward models that often fail to capture the mechanistic, nonlinear dose-response dynamics common in agronomy.
@@
In this work, we develop and analyze a new family of bandit algorithms that directly incorporate domain-specific nonlinear response models, offering greater interpretability and sample efficiency compared to classical linear or black-box approaches. Our approach is particularly valuable in settings where repeated decisions must be made with limited data, delayed feedback, and significant economic and environmental consequences.
Nonlinear models provide parsimonious and interpretable representations of domain-specific processes. Unlike generic polynomial or linear models, their parameters often map directly to biological quantities, such as maximum yield potential or nutrient efficiency, making the models both scientifically meaningful and statistically efficient. In agriculture, such models (Miguez et al.,, 2018) capture complex responses like crop yield to fertilizer, temperature effects on growth, or pest dynamics, where the relationship follows saturating, logistic, or exponential patterns rather than a simple linear trend. Their adoption is motivated not only by statistical accuracy but also by the need for transparent recommendations in sustainable resource management
@@ 1.1 Background and related work
Multi-armed bandit algorithms (Berry and Fristedt,, 1985; Lattimore and Szepesvári,, 2018) provide a core statistical framework for sequential decision-making under uncertainty, where the key challenge is to balance exploration and exploitation. In the original bandit formulation, each arm was associated with an unknown fixed reward distribution, and repeated pulls of the same arm were assumed to yield independent and identically distributed (i.i.d.) rewards. Subsequent work extended this framework to incorporate contextual information, leading to the development of contextual bandit models (Slivkins,, 2014; Chu et al.,, 2011). In these settings, regression models such as linear (Abbasi-Yadkori et al.,, 2011), generalized linear models (GLMs) (Li et al.,, 2010; Filippi et al.,, 2010), and more flexible nonparametric approaches, including kernels, nearest neighbors, and neural networks (Yang and Zhu,, 2002; Rigollet and Zeevi,, 2010; Srinivas et al.,, 2010; Valko et al.,, 2013), enable adaptive decision-making tailored to evolving arm/covariate information.
Classical nonlinear regression models, such as Mitscherlich (Dhanoa et al.,, 2022), Michaelis-Menten (Lopez et al.,, 2000), quadratic plateau models (Bélanger et al.,, 2000), and logistic response functions (Sepaskhah et al.,, 2011), are routinely used to capture crop yield response to fertilizer and other management variables. These models offer interpretable parameters tied to underlying biological processes, which is important for transparent and actionable recommendations (De Bruin and Pedersen,, 2008; Miguez et al.,, 2018).
@@
It is well-known that model-based bandit and reinforcement learning approaches can be far more sample-efficient than model-free methods when the underlying reward mechanisms are known (Osband and Van Roy,, 2014; Sun et al.,, 2019). However, such approaches have received limited attention in agricultural decision-making from a precision-agriculture and sustainability perspective, where established nonlinear models are often available but underutilized. Our work addresses this gap by integrating domain-driven nonlinear models directly into the bandit framework. This enables adaptive, interpretable, and sample-efficient resource management, advancing the practical and theoretical toolkit for sustainable and cost-effective fertilizer use.
@@
  • Theoretical guarantees: We provide regret and sample complexity results under both well-specified and shape-compatible misspecified models, and discuss their implications for agricultural decision-making under uncertainty.
  • Simulation studies emulating real-world practice: Through extensive simulations inspired by realistic fertilizer optimization scenarios, we demonstrate the practical benefits of our approach over classical linear and nonparametric bandit methods, particularly in low-sample regimes.
  • Interpretability and domain relevance: By leveraging mechanistic models with interpretable parameters, our methods yield actionable recommendations for practitioners, supporting transparent, resource-efficient, and economically sustainable decision-making.
We begin by setting up the problem in Section 2. In Section 3, we lay out the proposed framework for nonlinear model-based bandits, in which we study three model-based algorithms: $\epsilon$-greedy, Upper Confidence Bound (UCB), and the ViOlin algorithm. Section 4 presents known results for parametric bandits in both well-specified and misspecified settings, and derives sample complexity bounds for the specific nonlinear models considered by the ViOlin algorithm. We then conduct extensive simulation experiments emulating real-world yield production and fertilizer optimization scenarios, with the goal of resource-efficient and cost-effective management in sample-limited settings. Finally, Section 6 concludes the paper.
2 Problem Setup: Adaptive Input Decisions with Nonlinear Models
Formally, we consider a sequential decision-making scenario with a time horizon of $T$ rounds. Each round can represent, for example, a growing season or field trial, where decisions are made at planting and outcomes observed at harvest. At each round $t=1,\dots,T$, the decision-maker (agent) chooses an action (input level) $x_{t}\in\mathcal{X}$ from a feasible set $\mathcal{X}$, such as nitrogen application rates. After choosing action $x_{t}$, the agent observes a noisy reward $y_{t}$, representing yield or profit, modeled as:
@@ 2 Problem Setup: Adaptive Input Decisions with Nonlinear Models
This assumption reflects the variability of agricultural experiments driven by weather, soil heterogeneity, and biological processes, while ruling out extreme outliers that would make learning impractical.
The decision-maker seeks a policy/algorithm $\pi$ that selects actions based on historical observations to maximize cumulative expected reward. Performance can equivalently be measured through the cumulative regret, which quantifies the total loss relative to always applying the best treatment:
@@
where $x_{t}^{*}=\arg\max_{x\in\mathcal{X}}f(x;\beta^{*})$ denotes the optimal treatment at round $t$. In agricultural terms, regret measures the cumulative yield or profit lost due to not choosing the best possible input at each season. The objective is to devise algorithms achieving sublinear regret, meaning $R_{T}(\pi)=o_{P}(T)$, so that the average regret per round diminishes over time. Such a property guarantees that the policy rapidly identifies near-optimal input levels, minimizing wasted resources and experimental costs. Other performance metrics more suitable for risk-averse decision-making have also been proposed, such as the Conditional Value-at-Risk (CVaR) of the regret, which emphasizes control of the worst-case outcomes in the tail of the regret distribution (Baudry et al.,, 2021; Wang et al.,, 2023).
To illustrate the framework, we focus on fertilizer-yield relationships and investigate four classical nonlinear models widely used in agricultural research:
@@
    Michaelis-Menten Model: $f(x;a,b,d)=d+\frac{ax}{b+x}$, describing nutrient uptake or growth responses with diminishing returns.
@@
    Quadratic Plateau Model:
@@
    modeling responses that rise and then stabilize at a plateau, common in fertilizer trials.
@@
    Logistic Dose-Response Model: $f(x;A,B,C,d)=d+\frac{A}{1+\exp\left(-B(x-C)\right)}$, often used to capture threshold or inflection-point behavior.
These models are parsimonious, interpretable, and directly tied to biological processes. For example, in the Mitscherlich model, $A$ represents the maximum additional yield achievable with fertilizer, $b$ reflects the efficiency of nutrient uptake, and $d$ is the baseline yield. For example, a larger value of $b$ indicates that yield saturates more rapidly with increasing input, whereas a smaller $b$ corresponds to a more gradual response. Such parameters are routinely used in agronomic practice and make the resulting recommendations not only statistically grounded but also transparent and actionable.
Although we emphasize nitrogen fertilizer management as a motivating example, the framework generalizes to a wide range of agricultural and environmental problems where nonlinear dose-response relationships are well established, including irrigation efficiency, biomass accumulation, and nutrient loss in response to rainfall (Miguez et al.,, 2018). We provide detailed parameterizations and simulation examples in Section 5.
3 Methods: Algorithms for nonlinear model-based bandits
In agricultural trials, each season offers only limited opportunities for experimentation. Algorithms must therefore use past data efficiently while remaining robust to uncertainty. In this section, we describe a family of strategies for nonlinear model-based bandit problems, where the expected reward is modeled as a nonlinear function of the chosen input. We first provide a general algorithmic template, followed by detailed descriptions of three model-based strategies, namely, $\epsilon$-greedy, Upper Confidence Bound (UCB), and ViOlin, adapted to classical agronomic response models. Note that all the three algorithms have been studied in various bandit problems before (Auer and Ortner,, 2010; Arya and Yang,, 2020; Dong et al.,, 2021), here we specialize them to mechanistic nonlinear settings.
@@ 3.1 General framework
Algorithm 1 General Nonlinear Model-based Bandit Framework
@@
where $\ell$ is a loss function (such as squared-error loss) computed using the observed $(x_{i},y_{i})$ pairs, with the goal of minimizing the loss. This practical approach is central to our implementation, allowing bandit algorithms to leverage data-driven parameter estimates and domain-specific models. Then, in step 5, the decision-maker chooses the next action by balancing the exploration-exploitation trade-off and satisfying the criterion for cumulative reward maximization. Finally, in step 6, the outcome $y_{t}$ is observed, and is added to the dataset, enabling continual refinement of parameter estimates.
@@
Unlike classical supervised learning or factorial field experiments that provide outcomes for every treatment, bandit settings reveal only the reward for the action actually chosen at each round. This partial feedback, often referred to as bandit feedback, captures the reality of agricultural experimentation where only tested management options yield observable outcomes.
@@ 3.2 Epsilon-Greedy
The $\epsilon$-greedy algorithm (Algorithm 2) (Langford and Zhang,, 2008; Yang and Zhu,, 2002) is one of the simplest approaches yet effective approaches for balancing exploration and exploitation. At each round, the current nonlinear model is fitted using past observations, and the input that maximizes predicted profit is selected with high probability $1-\epsilon_{t}$ (exploitation). With small probability $\epsilon_{t}$, the algorithm explores by selecting a random alternative from the feasible input grid (exploration). This mechanism ensures continued learning, preventing the algorithm from becoming stuck on suboptimal choices. In agricultural contexts, this method reflects the intuitive practice of trying occasional alternative fertilizer rates even when one rate appears best.
@@
Algorithm 2 Nonlinear model-based Epsilon-Greedy